Find the upper and lower Riemann integrals of the function f , defined on [a,b]
as follows:

0 when is x irrational
f(x)={
2 when is x irrational

Is f Riemann integrable on [ a,b]?Justify your answer.
1
Expert's answer
2020-03-05T16:37:02-0500

Let's assume that "f=2" if "x\in Q"

For any partition "0<x_1<x_2<...x_n<1" the Riemann's upper integral is

"\lim\limits_{n\to\infty}\sum\limits_{k=1}^n(x_k-x_{k-1})\sup\limits_{x_{k-1}<x<x_k}f(x)=2\lim\limits_{n\to\infty}\sum\limits_{k=1}^n(x_k-x_{k-1})=2"

because in each interval "(x_{k-1};x_k)" there is a rational number "q" and hence the supremum is 2.

The lower integral is

"\lim\limits_{n\to\infty}\sum\limits_{k=1}^n(x_k-x_{k-1})\inf\limits_{x_{k-1}<x<x_k}f(x)=0"

because in each interval "(x_{k-1};x_k)" there is an irrational number "\alpha" and hence the infimum is 0.

Since upper and lower interals are not equal, the function is not Riemann integrable.